# mbrs: A Library for Minimum Bayes Risk Decoding

### v1

@@ E.2 Additional translation experiments on other language directions
QE kiwi 23.8 50.7 82.2 63.3 81.9 25.8 25.1 83.0 55.7 82.8 15.4 46.6 76.9 55.0 77.7 16.3 26.1 86.4 50.8 86.7 10.8 37.0 76.3 52.3 80.1
MBR bleu MC 27.2 52.1 78.5 62.0 75.3 29.8 27.8 77.3 53.8 74.2 17.2 46.8 71.9 51.8 69.4 17.9 26.2 81.5 48.2 80.0 11.7 35.5 70.1 47.4 69.9
MB 26.4 51.1 78.5 61.9 75.3 25.6 25.5 76.5 53.2 73.4 12.4 35.4 67.4 49.7 65.7 15.6 23.1 79.9 46.8 78.0 9.4 29.6 68.1 46.3 68.5
chrf MC 27.5 53.4 79.0 62.8 76.0 29.5 28.5 77.6 53.8 74.4 17.3 48.9 72.8 53.1 70.4 16.8 26.7 81.9 49.1 80.9 12.0 38.4 70.9 49.5 71.1
MB 26.8 51.9 78.8 62.5 76.0 25.4 25.8 76.8 53.4 73.7 12.3 35.9 67.9 50.5 66.2 15.6 23.4 80.1 47.3 78.6 9.8 31.0 68.8 47.7 69.5
comet MC 25.7 51.9 83.0 63.7 78.5 28.1 26.5 84.1 55.4 78.7 15.8 46.7 77.3 54.1 73.0 16.6 26.0 87.9 50.5 83.5 10.5 36.0 76.7 50.4 74.0
MB 26.2 51.4 81.6 63.2 77.6 25.5 25.7 81.4 54.5 76.6 12.3 35.6 69.9 50.8 67.4 15.7 23.6 84.0 48.7 80.5 9.6 30.4 72.2 48.4 71.2
bleurt MC OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
RAMBR bleu MC 27.3 52.1 78.6 62.1 75.4 29.8 27.7 77.3 53.7 74.2 17.2 46.8 72.1 52.0 69.6 17.7 26.0 81.5 48.0 80.0 11.8 35.6 70.1 47.8 70.0
MB 26.4 51.1 78.5 62.0 75.4 25.4 25.4 76.4 53.2 73.3 12.0 35.3 67.3 49.8 65.6 15.5 23.0 79.7 46.6 77.9 9.4 29.9 68.1 46.7 68.7
chrf MC 27.4 53.4 79.0 62.9 76.0 29.4 28.6 77.5 53.5 74.4 17.1 48.7 72.8 53.1 70.4 16.8 26.8 81.8 49.1 80.8 11.8 38.4 71.0 49.5 71.2
MB 26.9 51.9 78.9 62.6 76.0 25.1 25.7 76.6 53.0 73.6 12.2 35.8 67.8 50.5 66.2 15.3 23.4 80.0 47.3 78.6 9.8 31.1 68.8 47.9 69.5
comet MC 26.0 51.9 82.3 63.5 78.0 28.7 26.9 83.2 55.4 78.0 16.0 46.9 76.6 53.9 72.6 16.9 26.2 87.3 50.5 83.4 10.8 36.2 75.8 50.1 73.4
MB 26.2 51.4 81.0 63.1 77.2 25.6 25.7 80.5 54.3 76.0 12.3 35.7 69.6 50.7 67.2 15.9 23.6 83.3 48.3 80.1 9.4 30.3 71.5 48.1 70.6
CBMBR comet MC 25.1 51.4 83.0 63.3 78.3 27.6 26.0 84.0 55.2 78.5 15.7 46.5 77.2 53.7 72.8 16.5 26.0 87.8 50.1 83.3 10.0 35.4 76.6 50.1 73.5
MB 24.2 50.4 81.9 62.6 77.4 26.9 25.6 82.8 54.4 77.5 14.5 44.5 75.0 52.6 71.1 15.7 24.9 86.4 49.0 82.3 9.3 33.8 74.5 48.8 72.1
Prune- bleu MC 27.1 52.0 78.5 61.9 75.3 29.8 27.8 77.3 53.8 74.2 17.2 46.8 72.0 51.8 69.4 17.9 26.2 81.5 48.2 80.0 11.7 35.4 70.0 47.4 69.8
MBR MB 26.2 51.4 78.6 61.9 75.4 28.5 27.1 77.3 53.7 74.1 15.5 44.1 71.2 51.4 68.9 16.5 25.1 81.3 47.5 79.4 10.1 33.1 69.4 46.9 69.3
chrf MC 27.4 53.5 78.9 62.8 75.9 29.5 28.5 77.6 53.8 74.5 17.2 48.8 72.8 53.2 70.5 16.9 26.8 81.8 49.1 80.9 11.9 38.3 70.8 49.5 71.0
MB 26.8 52.3 78.8 62.4 75.9 28.8 27.4 77.6 53.9 74.4 15.9 45.2 71.7 52.2 69.4 16.6 25.4 81.5 48.0 79.9 10.8 34.9 70.2 48.4 70.3
comet MC 25.8 51.9 83.0 63.7 78.4 28.1 26.5 84.1 55.5 78.6 15.9 46.8 77.3 54.1 72.9 16.5 26.0 87.9 50.4 83.6 10.4 35.9 76.6 50.3 73.9
MB 26.3 51.9 81.6 63.3 77.6 28.9 27.3 82.3 55.1 77.4 15.2 44.3 74.4 52.7 70.9 17.0 25.7 86.1 49.7 82.3 10.0 34.1 73.9 49.2 72.2
bleurt MC 24.9 51.3 79.4 65.1 76.4 27.3 25.9 78.0 58.7 75.0 15.7 46.5 73.6 56.7 71.3 15.8 25.5 82.5 53.9 80.6 10.2 35.2 71.3 53.2 71.4
MB 26.0 51.5 79.1 63.5 76.1 28.3 26.9 77.7 55.7 74.6 15.1 44.2 72.0 53.7 69.7 16.2 25.0 81.9 50.0 80.0 9.9 33.2 70.4 50.0 70.4
PMBR bleu MC 26.4 51.4 78.2 61.8 75.0 29.0 27.1 77.0 53.7 73.9 16.9 46.5 71.8 51.7 69.2 17.4 25.7 81.1 48.1 79.7 11.2 34.9 69.9 47.4 69.8
MB 26.4 51.5 78.0 61.6 74.7 28.9 27.0 76.7 53.2 73.7 16.5 45.8 71.4 51.4 68.8 17.3 25.7 81.2 47.8 79.6 11.1 34.6 69.4 47.1 69.2
chrf MC 26.0 52.2 78.3 62.1 75.1 29.0 27.9 77.3 53.7 74.2 16.0 47.6 72.2 52.5 69.8 16.4 26.1 81.3 48.5 80.2 11.1 37.2 70.4 49.0 70.6
MB 25.7 52.1 78.1 61.9 75.0 28.7 27.8 76.9 53.1 73.8 15.9 47.4 72.0 52.4 69.6 16.3 26.0 81.1 48.2 79.9 11.2 37.2 70.1 48.7 70.2
comet MC 25.9 51.9 82.8 63.6 78.3 28.3 26.6 83.9 55.5 78.5 15.9 46.7 77.1 54.0 72.9 16.7 26.0 87.7 50.3 83.4 10.5 35.8 76.5 50.3 73.8
MB 25.7 51.5 82.1 63.5 77.9 27.0 26.1 82.5 54.9 77.4 14.1 41.5 73.5 52.3 70.0 16.0 24.7 85.9 49.4 82.1 10.0 32.6 74.0 49.1 72.1
bleurt MC 25.0 51.4 79.4 65.1 76.4 27.6 26.2 78.0 58.6 74.9 15.5 46.4 73.5 56.6 71.2 15.8 25.5 82.5 53.7 80.6 10.1 35.2 71.5 53.2 71.6
MB 25.5 51.4 79.2 64.1 76.3 27.7 26.3 77.6 56.9 74.5 14.7 44.1 72.0 54.4 69.6 16.0 24.7 82.0 51.2 79.9 9.6 33.5 70.5 51.1 70.7
Oracle bleu 46.4 63.4 80.9 66.6 74.3 45.5 41.9 80.9 61.4 74.2 31.9 54.9 74.4 56.3 69.0 34.7 39.0 83.4 54.4 78.7 28.0 46.2 73.2 53.1 69.5
chrf 43.5 66.0 81.3 67.4 74.8 45.6 42.8 81.0 61.3 74.1 28.2 58.2 75.5 57.8 70.0 33.2 41.4 84.2 56.3 79.2 24.0 50.4 74.5 55.2 70.8
comet 34.4 58.3 86.4 68.5 78.9 35.4 33.4 87.6 61.8 79.3 20.5 50.8 81.2 58.9 73.9 22.4 31.6 90.6 56.8 83.6 15.9 41.2 81.9 56.3 75.5
bleurt 35.7 59.0 82.5 72.6 76.3 36.3 34.2 81.9 68.6 75.6 21.3 50.9 76.6 63.8 71.6 23.5 33.0 85.3 64.3 80.4 16.3 41.6 75.9 62.2 72.8

### v2

@@ D.2 Additional translation experiments on other language directions
En–Ja Ja–En
Decoding Utility Est. Bleu chrF Comet Bleurt Kiwi Bleu chrF Comet Bleurt Kiwi
MAPϵ 15.1 22.6 79.2 46.1 77.4 9.0 29.2 68.0 46.5 68.5
@@
QE Kiwi 16.3 26.1 86.4 50.8 86.7 10.8 37.0 76.3 52.3 80.1
MBR Bleu MC 17.9 26.2 81.5 48.2 80.0 11.7 35.5 70.1 47.4 69.9
MB 15.6 23.1 79.9 46.8 78.0 9.4 29.6 68.1 46.3 68.5
chrF MC 16.8 26.7 81.9 49.1 80.9 12.0 38.4 70.9 49.5 71.1
MB 15.6 23.4 80.1 47.3 78.6 9.8 31.0 68.8 47.7 69.5
Comet MC 16.6 26.0 87.9 50.5 83.5 10.5 36.0 76.7 50.4 74.0
MB 15.7 23.6 84.0 48.7 80.5 9.6 30.4 72.2 48.4 71.2
Bleurt MC OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
MB OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
RAMBR Bleu MC 17.7 26.0 81.5 48.0 80.0 11.8 35.6 70.1 47.8 70.0
MB 15.5 23.0 79.7 46.6 77.9 9.4 29.9 68.1 46.7 68.7
chrF MC 16.8 26.8 81.8 49.1 80.8 11.8 38.4 71.0 49.5 71.2
MB 15.3 23.4 80.0 47.3 78.6 9.8 31.1 68.8 47.9 69.5
Comet MC 16.9 26.2 87.3 50.5 83.4 10.8 36.2 75.8 50.1 73.4
MB 15.9 23.6 83.3 48.3 80.1 9.4 30.3 71.5 48.1 70.6
CBMBR Comet MC 16.5 26.0 87.8 50.1 83.3 10.0 35.4 76.6 50.1 73.5
MB 15.7 24.9 86.4 49.0 82.3 9.3 33.8 74.5 48.8 72.1
Pruning Bleu MC 17.9 26.2 81.5 48.2 80.0 11.7 35.4 70.0 47.4 69.8
MB 16.5 25.1 81.3 47.5 79.4 10.1 33.1 69.4 46.9 69.3
chrF MC 16.9 26.8 81.8 49.1 80.9 11.9 38.3 70.8 49.5 71.0
MB 16.6 25.4 81.5 48.0 79.9 10.8 34.9 70.2 48.4 70.3
Comet MC 16.5 26.0 87.9 50.4 83.6 10.4 35.9 76.6 50.3 73.9
MB 17.0 25.7 86.1 49.7 82.3 10.0 34.1 73.9 49.2 72.2
Bleurt MC 15.8 25.5 82.5 53.9 80.6 10.2 35.2 71.3 53.2 71.4
MB 16.2 25.0 81.9 50.0 80.0 9.9 33.2 70.4 50.0 70.4
PMBR Bleu MC 17.4 25.7 81.1 48.1 79.7 11.2 34.9 69.9 47.4 69.8
MB 17.3 25.7 81.2 47.8 79.6 11.1 34.6 69.4 47.1 69.2
chrF MC 16.4 26.1 81.3 48.5 80.2 11.1 37.2 70.4 49.0 70.6
MB 16.3 26.0 81.1 48.2 79.9 11.2 37.2 70.1 48.7 70.2
Comet MC 16.7 26.0 87.7 50.3 83.4 10.5 35.8 76.5 50.3 73.8
MB 16.0 24.7 85.9 49.4 82.1 10.0 32.6 74.0 49.1 72.1
Bleurt MC 15.8 25.5 82.5 53.7 80.6 10.1 35.2 71.5 53.2 71.6
MB 16.0 24.7 82.0 51.2 79.9 9.6 33.5 70.5 51.1 70.7
Oracle Bleu 34.7 39.0 83.4 54.4 78.7 28.0 46.2 73.2 53.1 69.5
chrF 33.2 41.4 84.2 56.3 79.2 24.0 50.4 74.5 55.2 70.8
Comet 22.4 31.6 90.6 56.8 83.6 15.9 41.2 81.9 56.3 75.5
Bleurt 23.5 33.0 85.3 64.3 80.4 16.3 41.6 75.9 62.2 72.8
[…]